# Massless Quasiparticles in Bogoliubov-de Gennes Systems

### v1

@@ E.1 When the $\tilde{\mathcal{P}}$ is unitary (chiral-like, such as $\tilde{\mathcal{P}}=\mathcal{TP}$)
From $D(\tilde{\mathcal{P}})=MD^{\prime}(\tilde{\mathcal{P}})$ one has
$[D(\tilde{\mathcal{P}})]^{2}=MD^{\prime}(\tilde{\mathcal{P}})MD^{\prime}(\tilde{\mathcal{P}}).$
@@
From the above two relations, we have $D^{\prime}(\tilde{\mathcal{P}})=MD^{\prime}(\tilde{\mathcal{P}})M$, or equivalently
$D(\tilde{\mathcal{P}})=MD(\tilde{\mathcal{P}})M.$ (42)
@@
If $G_{f}(k)$ is unitary, the unitary elements in the centralizer of $D(G_{f}(k))$ form a $U(1)$ group $\{e^{\mathrm{i}\theta}I;\theta\in[0,2\pi)\}$ according to Schur’s lemma. Since $M$ belongs to the unitary centralizer, we have $M=\eta_{\tilde{\mathcal{P}}}I$ with $|\eta_{\tilde{\mathcal{P}}}|=1$, namely $D(\tilde{\mathcal{P}})=\eta_{\tilde{\mathcal{P}}}D^{\prime}(\tilde{\mathcal{P}})$.
@@
so there exists a nonzero hermitian Hamiltonian $h_{k}=\varepsilon_{k}\mu_{x}\otimes I$ commuting with $\tilde{D}(G_{f}(k))$ and anti-commuting with $\tilde{D}(\tilde{\mathcal{P}})$. Namely, this case corresponds to reducible minimal nonzero modes.
@@
so there also exists a nonzero Hamiltonian $h_{k}=\varepsilon_{k}\mu_{x}\otimes I$ commuting with $\tilde{D}(G_{f}(k))$ and anticommuting with $\tilde{D}(\tilde{\mathcal{P}})$. This gives rise to a set of reducible minimal nonzero modes.
(b2) When $D(G_{f}(k))$ is of the complex class $\mathbb{C}$, the $D(H_{f}(k))$ is a direct sum of two non-equivalent irreps. The centralizer of $D(G_{f}(k))$ takes the form $\left\{\begin{pmatrix}aI&0\\ 0&a^{*}I\end{pmatrix},a\in\mathbb{C}\right\}$. Hence $M$ belongs to the $U(1)$ group $M\in\left\{e^{\mathrm{i}\theta\Omega_{z}},\theta\in[0,2\pi)\right\}$ formed by unitary elements in the centralizer.
Letting $M=e^{i\theta_{0}\Omega_{z}}$, since $M$ falls in the centralizer of $D(G_{f}(k)$, for any $g\in G_{f}(k)$, one has
@@
Namely, under the transformation $e^{-\mathrm{i}\Omega_{z}{\theta_{0}\over 2}}$, $D(\tilde{\mathcal{P}})$ is transformed into $D^{\prime}(\tilde{\mathcal{P}})=e^{-\mathrm{i}\theta_{0}\Omega_{z}}D(\tilde{\mathcal{P}})$ without affecting the reps of elements in $G_{f}(k)$. In the case $\theta_{0}=0$, the nonzero hermitian hamiltonian can be chosen as $h_{k}=\varepsilon_{k}\begin{pmatrix}0&-\mathrm{i}\Omega_{z}\\ \mathrm{i}\Omega_{z}&0\end{pmatrix}=\varepsilon_{k}\mu_{y}\otimes\Omega_{z}$. The Hamiltonian with nonzero $\theta_{0}$ can be obtained with the transformation (47) of the second rep. This gives rise to a set of reducible minimal nonzero modes.
@@
there exists a nonzero hermitian Hamiltonian $h_{k}=\varepsilon_{k}\mu_{x}\otimes I$. This gives rise to a set of reducible minimal nonzero modes.
(b3) When $D(G_{f}(k))$ is of the quaternionic class $\mathbb{H}$, the rep $D(H_{f}(k))$ is a direct sum of two equivalent irreps. Generally, the centralizer of $D(G_{f}(k))$ is $\left\{\begin{pmatrix}aI&bI\\ -b^{*}I&a^{*}I\end{pmatrix},a,b\in\mathbb{C}\right\}$, whose unitary elements form an $SU(2)$ group $\left\{\begin{pmatrix}aI&bI\\ -b^{*}I&a^{*}I\end{pmatrix},aa^{*}+bb^{*}=1,a,b\in\mathbb{C}\right\}$. Since the unitary matrix $M=\left[D(\tilde{\mathcal{P}})D^{\prime-1}(\tilde{\mathcal{P}})\right]$ falls in the centralizer of $D(G_{f}(k))$, we can write
$M=\cos\theta_{0}+\sin\theta_{0}\left[n_{x}(\mathrm{i}\Omega_{x})+n_{y}(\mathrm{i}\Omega_{y})+n_{z}(\mathrm{i}\Omega_{z})\right]$
@@
$M=e^{\mathrm{i}\theta_{0}\Omega_{n}}=\cos\Phi I+\mathrm{i}\sin\theta_{0}\Omega_{n}.$ (53)
@@
for certain $\boldsymbol{n}_{0}\in\bm{S}^{2}$, then any value $\theta_{0}\in[0,2\pi)$ in the above $\boldsymbol{n}_{0}$ direction satisfies (54). The discussions in the $\mathbb{C}$-class also applies here. In the case $\theta_{0}=0$, one possible Hamiltonian is $h_{k}=\varepsilon_{k}\begin{pmatrix}0&\mathrm{i}\Omega_{z}\\ -\mathrm{i}\Omega_{z}&0\end{pmatrix}=\varepsilon_{k}\mu_{y}\otimes\Omega_{z}$. This gives rise to a set of reducible minimal nonzero modes.
@@
When $G_{f}(k)$ is unitary, then the unitary elements in the centralizer of $D(G_{f}(k))$ form a $U(1)$ group $\{e^{\mathrm{i}\theta}I;\theta\in[0,2\pi)\}$. Lemma1 indicates that $M$ belongs to the above $U(1)$ group. Letting $M=e^{\mathrm{i}\theta_{0}}$, then $D^{\prime}(\tilde{\mathcal{P}})K=e^{-\mathrm{i}\theta_{0}}D(\tilde{\mathcal{P}})K$ can be transformed into $D(\tilde{\mathcal{P}})K$ by adjusting the phase of the bases. Hence $D^{\prime}(F_{f}(k))$ is always equivalent to $D(F_{f}(k))$. Then there always exists a Hamiltonian $h_{k}=\varepsilon_{k}(\sin{\theta_{0}\over 2}\sigma_{x}+\cos{\theta_{0}\over 2}\sigma_{y})\otimes I_{n}$ which commutes with $\tilde{D}(h)=D(h)\oplus D(h)$ and anti-commutes with $\tilde{D}(\tilde{\mathcal{P}})K=D(\tilde{\mathcal{P}})\oplus e^{-\mathrm{i}\theta_{0}}D(\tilde{\mathcal{P}})K$. This gives rise to a set of reducible minimal nonzero modes.
@@ F.1 Proof of the theorem 2
$M(G_{f}(k))=m_{1}(G_{f}(k))\oplus m_{2}(G_{f}(k))\oplus...$
under the bases $[d_{k1}^{\dagger},d_{k2}^{\dagger},...,d_{kN}^{\dagger}]=[c_{k1}^{\dagger},c_{k2}^{\dagger},...,c_{kN}^{\dagger}]U$. Since $\tilde{\mathcal{P}}$ transforms creation operators into annihilation operators, we denote $[\tilde{d}_{k1},\tilde{d}_{k2},...,\tilde{d}_{kN}]=\widehat{\tilde{\mathcal{P}}}[d_{k1}^{\dagger},d_{k2}^{\dagger},...,d_{kN}^{\dagger}]\widehat{\tilde{\mathcal{P}}}^{-1}$. Since $\tilde{\mathcal{P}}G_{f}(k)=G_{f}(k)\tilde{\mathcal{P}}$, $[\tilde{d}_{k1},\tilde{d}_{k2},...,\tilde{d}_{kN}]$ carry a rep of $G_{f}(k)$,
$\tilde{M}(G_{f}(k))=\tilde{m}_{1}(G_{f}(k))\oplus\tilde{m}_{2}(G_{f}(k))\oplus...$
with $\tilde{m}_{a}(g)={\omega_{2}(g,\tilde{\mathcal{P}})\over\omega_{2}(\tilde{\mathcal{P}},\tilde{\mathcal{P}}^{-1}g\tilde{\mathcal{P}})}K_{s(\tilde{\mathcal{P}})}m_{a}(\tilde{\mathcal{P}}^{-1}g\tilde{\mathcal{P}})K_{s(\tilde{\mathcal{P}})}$ for $a=1,2,...$ and $g\in G_{f}(k)$.
In the direct sum space of $m_{a}(G_{f}(k))\oplus\tilde{m}_{a}(G_{f}(k))$, the group $F_{f}(k)=G_{f}(k)+\tilde{\mathcal{P}}G_{f}(k)$ is presented as
$D_{a}(g)K_{s(g)}=\begin{pmatrix}m_{a}(g)&0\\ 0&\tilde{m}_{a}(g)\end{pmatrix}K_{s(g)},\ \ \ g\in G_{f}(k)$
$D_{a}(\tilde{\mathcal{P}})K_{s(\tilde{\mathcal{P}})}=\left(\begin{matrix}0&\omega_{2}(\tilde{\mathcal{P}},\tilde{\mathcal{P}})m_{a}(\tilde{\mathcal{P}}^{2})\\ I_{n_{a}}&0\end{matrix}\right)K_{s(\tilde{\mathcal{P}})},$
with $n_{a}$ the dimensionality of the rep $m_{a}(G_{f}(k))$. In this space, there exists a non-degenerate Hamiltonian $h_{k}=\varepsilon_{k}\mu_{z}\otimes I_{n_{a}}$ with $\varepsilon_{k}\neq 0$ lifting all the zero modes. Notice that the above rep of $F_{f}(k)$ may be reducible if $m_{a}(g)=\tilde{m}_{a}(g)$ for all $g\in G_{f}(k)$. Therefore, in electron bases the nonzero energy modes have two possible resources:
(I) an irrep $D_{a}(F_{f}(k))$, in which the restricted rep $D_{a}(G_{f}(k))=m_{a}(G_{f}(k))\oplus\tilde{m}_{a}(G_{f}(k))$ is reducible and $m_{a}\neq\tilde{m}_{a}$, give rise to a set of irreducible nonzero modes;
(II) a direct sum of two irreps $D_{a}(F_{f}(k))=\mathcal{D}_{+}(F_{f}(k))\oplus\mathcal{D}_{-}(F_{f}(k))$, in which $\mathcal{D}_{+}(G_{f}(k))=m_{a}(G_{f}(k))=\mathcal{D}_{-}(G_{f}(k))=\tilde{m}_{a}(G_{f}(k))$, couple with each other to form a set of reducible minimal nonzero modes.
The above theorem implies that any set of irreducible zero modes contained in representations of $F_{f}(k)$ for arbitrary $\boldsymbol{k}$ can find its coupling partner (namely there only exist minimal nonzero modes) if the $F_{f}(k)$ doesn’t contain any ‘non-diagonal’ (effective) charge conjugation symmetry.
@@ F.2 Extension and induction: from the rep of a normal subgroup $H$ to the full group $G\rhd H$
as the complete little co-group at $\boldsymbol{k}$, then we have $F_{f}^{c}(k)/F_{f}(k)\cong{Z}^{c}_{2}$. Similarly, we can define $G_{f}^{c}(k)$ with $G_{f}^{c}(k)=G_{f}(k)+\mathcal{C}G_{f}(k)$. In the presence of $\mathcal{C}$, a set of minimal nonzero modes can be turned into zero modes in the following two ways:
(1) Two sets of irreducible zero modes contained in reducible minimal nonzero modes which are previously coupled becomes uncoupleable;
(2) A set of irreducible nonzero modes or reducible minimal nonzero modes are turned into a set of irreducible zero mode which has no coupling partner.

### v3

@@ E.1 When the $\tilde{\mathcal{P}}$ is unitary (chiral-like, such as $\tilde{\mathcal{P}}=\mathcal{TP}$)
From $D(\tilde{\mathcal{P}})=XD^{\prime}(\tilde{\mathcal{P}})$ one has
$[D(\tilde{\mathcal{P}})]^{2}=XD^{\prime}(\tilde{\mathcal{P}})XD^{\prime}(\tilde{\mathcal{P}}).$
@@
From the above two relations, we have $D^{\prime}(\tilde{\mathcal{P}})=XD^{\prime}(\tilde{\mathcal{P}})X$, or equivalently
$D(\tilde{\mathcal{P}})=XD(\tilde{\mathcal{P}})X.$ (42)
@@
If $G_{f}(k)$ is unitary, the unitary elements in the centralizer of $D(G_{f}(k))$ form a $U(1)$ group $\{e^{\mathrm{i}\theta}I;\theta\in[0,2\pi)\}$ according to Schur’s lemma. Since $X$ belongs to the unitary centralizer, we have $X=\eta_{\tilde{\mathcal{P}}}I$ with $|\eta_{\tilde{\mathcal{P}}}|=1$, namely $D(\tilde{\mathcal{P}})=\eta_{\tilde{\mathcal{P}}}D^{\prime}(\tilde{\mathcal{P}})$.
@@
so there exists a nonzero hermitian Hamiltonian $h_{k}=\varepsilon_{k}\mu_{x}\otimes I$ commuting with $\tilde{D}(G_{f}(k))$ and anti-commuting with $\tilde{D}(\tilde{\mathcal{P}})$. Namely, this case corresponds to RMNZM.
@@
so there also exists a nonzero Hamiltonian $h_{k}=\varepsilon_{k}\mu_{x}\otimes I$ commuting with $\tilde{D}(G_{f}(k))$ and anticommuting with $\tilde{D}(\tilde{\mathcal{P}})$. This gives rise to a set of RMNZM.
(b2) When $D(G_{f}(k))$ is of the complex class $\mathbb{C}$, the $D(H_{f}(k))$ is a direct sum of two non-equivalent irreps. The centralizer of $D(G_{f}(k))$ takes the form $\left\{\begin{pmatrix}aI&0\\ 0&a^{*}I\end{pmatrix},a\in\mathbb{C}\right\}$. Hence $X$ belongs to the $U(1)$ group $X\in\left\{e^{\mathrm{i}\theta\Omega_{z}},\theta\in[0,2\pi)\right\}$ formed by unitary elements in the centralizer.
Letting $X=e^{i\theta_{0}\Omega_{z}}$, since $X$ falls in the centralizer of $D(G_{f}(k)$, for any $g\in G_{f}(k)$, one has
@@
Namely, under the transformation $e^{-\mathrm{i}\Omega_{z}{\theta_{0}\over 2}}$, $D(\tilde{\mathcal{P}})$ is transformed into $D^{\prime}(\tilde{\mathcal{P}})=e^{-\mathrm{i}\theta_{0}\Omega_{z}}D(\tilde{\mathcal{P}})$ without affecting the reps of elements in $G_{f}(k)$. In the case $\theta_{0}=0$, the nonzero hermitian hamiltonian can be chosen as $h_{k}=\varepsilon_{k}\begin{pmatrix}0&-\mathrm{i}\Omega_{z}\\ \mathrm{i}\Omega_{z}&0\end{pmatrix}=\varepsilon_{k}\mu_{y}\otimes\Omega_{z}$. The Hamiltonian with nonzero $\theta_{0}$ can be obtained with the transformation (47) of the second rep. This gives rise to a set of RMNZM.
@@
there exists a nonzero hermitian Hamiltonian $h_{k}=\varepsilon_{k}\mu_{x}\otimes I$. This gives rise to a set of RMNZM.
(b3) When $D(G_{f}(k))$ is of the quaternionic class $\mathbb{H}$, the rep $D(H_{f}(k))$ is a direct sum of two equivalent irreps. Generally, the centralizer of $D(G_{f}(k))$ is $\left\{\begin{pmatrix}aI&bI\\ -b^{*}I&a^{*}I\end{pmatrix},a,b\in\mathbb{C}\right\}$, whose unitary elements form an $SU(2)$ group $\left\{\begin{pmatrix}aI&bI\\ -b^{*}I&a^{*}I\end{pmatrix},aa^{*}+bb^{*}=1,a,b\in\mathbb{C}\right\}$. Since the unitary matrix $X=\left[D(\tilde{\mathcal{P}})D^{\prime-1}(\tilde{\mathcal{P}})\right]$ falls in the centralizer of $D(G_{f}(k))$, we can write
$X=\cos\theta_{0}+\sin\theta_{0}\left[n_{x}(\mathrm{i}\Omega_{x})+n_{y}(\mathrm{i}\Omega_{y})+n_{z}(\mathrm{i}\Omega_{z})\right]$
@@
$X=e^{\mathrm{i}\theta_{0}\Omega_{n}}=\cos\Phi I+\mathrm{i}\sin\theta_{0}\Omega_{n}.$ (53)
@@
for certain $\boldsymbol{n}_{0}\in\bm{S}^{2}$, then any value $\theta_{0}\in[0,2\pi)$ in the above $\boldsymbol{n}_{0}$ direction satisfies (54). The discussions in the $\mathbb{C}$-class also applies here. In the case $\theta_{0}=0$, one possible Hamiltonian is $h_{k}=\varepsilon_{k}\begin{pmatrix}0&\mathrm{i}\Omega_{z}\\ -\mathrm{i}\Omega_{z}&0\end{pmatrix}=\varepsilon_{k}\mu_{y}\otimes\Omega_{z}$. This gives rise to a set of RMNZM.
@@
When $G_{f}(k)$ is unitary, then the unitary elements in the centralizer of $D(G_{f}(k))$ form a $U(1)$ group $\{e^{\mathrm{i}\theta}I;\theta\in[0,2\pi)\}$. Lemma1 indicates that $X$ belongs to the above $U(1)$ group. Letting $X=e^{\mathrm{i}\theta_{0}}$, then $D^{\prime}(\tilde{\mathcal{P}})K=e^{-\mathrm{i}\theta_{0}}D(\tilde{\mathcal{P}})K$ can be transformed into $D(\tilde{\mathcal{P}})K$ by adjusting the phase of the bases. Hence $D^{\prime}(F_{f}(k))$ is always equivalent to $D(F_{f}(k))$. Then there always exists a Hamiltonian $h_{k}=\varepsilon_{k}(\sin{\theta_{0}\over 2}\sigma_{x}+\cos{\theta_{0}\over 2}\sigma_{y})\otimes I_{n}$ which commutes with $\tilde{D}(h)=D(h)\oplus D(h)$ and anti-commutes with $\tilde{D}(\tilde{\mathcal{P}})K=D(\tilde{\mathcal{P}})\oplus e^{-\mathrm{i}\theta_{0}}D(\tilde{\mathcal{P}})K$. This gives rise to a set of RMNZM.
@@ F.1 Proof of the theorem 2
$M(G_{f}(k))=d_{1}(G_{f}(k))\oplus d_{2}(G_{f}(k))\oplus...$
under the bases $[b_{k1}^{\dagger},b_{k2}^{\dagger},...,b_{kN}^{\dagger}]=[c_{k1}^{\dagger},c_{k2}^{\dagger},...,c_{kN}^{\dagger}]U$. Since $\tilde{\mathcal{P}}$ transforms creation operators into annihilation operators, we denote $[\tilde{b}_{k1},\tilde{b}_{k2},...,\tilde{b}_{kN}]=\widehat{\tilde{\mathcal{P}}}[b_{k1}^{\dagger},b_{k2}^{\dagger},...,b_{kN}^{\dagger}]\widehat{\tilde{\mathcal{P}}}^{-1}$. Since $\tilde{\mathcal{P}}G_{f}(k)=G_{f}(k)\tilde{\mathcal{P}}$, $[\tilde{b}_{k1},\tilde{b}_{k2},...,\tilde{b}_{kN}]$ carry a rep of $G_{f}(k)$,
$\tilde{M}(G_{f}(k))=\tilde{d}_{1}(G_{f}(k))\oplus\tilde{d}_{2}(G_{f}(k))\oplus...$
with $\tilde{d}_{a}(g)={\omega_{2}(g,\tilde{\mathcal{P}})\over\omega_{2}(\tilde{\mathcal{P}},\tilde{\mathcal{P}}^{-1}g\tilde{\mathcal{P}})}K_{s(\tilde{\mathcal{P}})}d_{a}(\tilde{\mathcal{P}}^{-1}g\tilde{\mathcal{P}})K_{s(\tilde{\mathcal{P}})}$ for $a=1,2,...$ and $g\in G_{f}(k)$.
In the direct sum space of $d_{a}(G_{f}(k))\oplus\tilde{d}_{a}(G_{f}(k))$, the group $F_{f}(k)=G_{f}(k)+\tilde{\mathcal{P}}G_{f}(k)$ is presented as
$\!\!\!\!\!\!\!\!\!\!D_{a}(g)K_{s(g)}=\begin{pmatrix}d_{a}(g)&0\\ 0&\tilde{d}_{a}(g)\end{pmatrix}K_{s(g)},\ \ \ g\in G_{f}(k),$ (55)
$\!\!\!\!\!\!\!\!\!\!D_{a}(\tilde{\mathcal{P}})K_{s(\tilde{\mathcal{P}})}=\left(\begin{matrix}0&\omega_{2}(\tilde{\mathcal{P}},\tilde{\mathcal{P}})d_{a}(\tilde{\mathcal{P}}^{2})\\ I_{n_{a}}&0\end{matrix}\right)K_{s(\tilde{\mathcal{P}})},$ (56)
with $n_{a}$ the dimensionality of the rep $d_{a}(G_{f}(k))$. In this space, there exists a non-degenerate Hamiltonian
$h_{k}=\varepsilon_{k}\mu_{z}\otimes I_{n_{a}}$
with $\varepsilon_{k}\neq 0$ lifting all the zero modes. Notice that the above rep (55) and (56) of $F_{f}(k)$ may be reducible if $d_{a}(g)=\tilde{d}_{a}(g)$ for all $g\in G_{f}(k)$. Therefore, in electron bases the nonzero energy modes have two possible resources:
(I) a composite irrep $D_{a}(F_{f}(k))$, in which the restricted rep $D_{a}(G_{f}(k))=d_{a}(G_{f}(k))\oplus\tilde{d}_{a}(G_{f}(k))$ is reducible and $d_{a}\neq\tilde{d}_{a}$, give rise to a set of irreducible nonzero modes;
(II) a direct sum of two simple irreps $D_{a}(F_{f}(k))=\mathcal{D}_{+}(F_{f}(k))\oplus\mathcal{D}_{-}(F_{f}(k))$, in which $\mathcal{D}_{+}(G_{f}(k))=d_{a}(G_{f}(k))=\mathcal{D}_{-}(G_{f}(k))=\tilde{d}_{a}(G_{f}(k))$, couple with each other to form a set of reducible minimal nonzero modes.
The above theorem implies that any simple irrep of $F_{f}(k)$ at arbitrary $\boldsymbol{k}$ can find its coupling partner to form nonzero energy modes. Hence we conclude that without additional ‘non-diagonal’ symmetries like the (effective) charge conjugation symmetry, zero energy modes are unstable at any given $\boldsymbol{k}$.
@@ F.2 Extension and induction: from the rep of a normal subgroup $H$ to the full group $G\rhd H$
as the complete little co-group at $\boldsymbol{k}$, then we have $F_{f}^{c}(k)/F_{f}(k)\cong{Z}_{2}$. Similarly, we can define $G_{f}^{c}(k)$ with $G_{f}^{c}(k)=G_{f}(k)+\mathcal{C}G_{f}(k)$ and $G^{c}_{f}(k)/G_{f}(k)\cong{Z}_{2}$. In the presence of $\mathcal{C}$, a set of minimal nonzero modes can be turned into zero modes (called $\mathcal{C}$-zero modes) in the following two ways:
(1) Two simple irreps contained in RMNZM which are previously coupled become uncoupleable and form reducible minimal $\mathcal{C}$-zero modes;
(2) A set of irreducible nonzero modes or a set of RMNZM are turned into a set of irreducible $\mathcal{C}$-zero modes.